OpenStudy (anonymous):

y = -1/3x + 3 –x + 3y = 9

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

i need to know the y-intercept of the equation.

OpenStudy (diyadiya):

Oops sorry ignore my post :/

OpenStudy (anonymous):

and the slope

OpenStudy (diyadiya):

Which is the equation?

OpenStudy (anonymous):

y = -1/3x + 3 i think this one is the y-intercept

OpenStudy (anonymous):

–x + 3y = 9 and this is the slope

OpenStudy (diyadiya):

Sorry, do you have two equations? y-intercept is when x=0 to find the slope write the equation in the form y=mx + b where 'm' is the slope & b is the y-intercept

OpenStudy (anonymous):

ok thank you

OpenStudy (diyadiya):

Welcome :)